91914108, 16) the total current (a) equivalent resistaice, (b) total in the following circuit, (c) the current through each e drop across each resistor, and (e) the power used by each. Showing work resistor resistor, (d) the voltage drop 2. 2o R0 4o 96 40 12 13 20 I0 Ri R2 R3 Pi- P2= 1 -40
Answer #1

Solution :

a)All three resistors are in series so there equivalent resistance is

R_{eq}=R_1+R_2+R_3 = 8+10+2=20\Omega

-------------------------------------------------------------------

b)

Total current in circuit is given by

I = \frac{V}{R_{eq}} = \frac{40}{20}=2 A

-------------------------------------------------------------

c)

All resistors are in series so all have same current through them = I = 2 A.

---------------------------------------------------------------------

d)

Voltage drop across each can be found using ohm's law

1.Across R1 = V1 = I R1 = 2 \times 8 = 16 V

2..Across R2 = V2 = I R2 = 2 \times 10 = 20 V

3..Across R3 = V3 = I R3 = 2 \times 2 = 4 V

--------------------------------------------

e)

Power (P) used by each can be calculated using , P = VI

For R1 , P1 = V1I = 16 \times 2 = 32 W

For R2 , P2 = V2I = 20 \times 2 = 40 W

For R3 , P3 = V3I = 4\times 2 = 8 W